Coronado Company constructed a building at a cost of $2,464,000 and occupied it beginning in January 1998. It was estimated at that time that its life would be 40 years, with no salvage value.

In January 2018, a new roof was installed at a cost of $336,000, and it was estimated then that the building would have a useful life of 25 years from that date. The cost of the old roof was $179,200.

What amount of depreciation should have been charged annually from the years 2001 to 2020? (Assume straight-line depreciation.)

Depreciation from the years 2001 to 2020

$

What entry should be made in 2021 to record the replacement of the roof?

Prepare the entry in January 2021 to record the revision in the estimated life of the building, if necessary.

What amount of depreciation should be charged for the year 2021?

Depreciation for the year 2021 (Assume the cost of the old roof is removed)

$

Depreciation for the year 2021 (Assume the cost of the new roof is debited to accumulated depreciation - building)
Answer #1

Ans: A). Depreciation for the year 2001 to 2020:cost- salvage value/ years

=(2464000-0)/40

=$61,600

B). To record replacement of the roof:

S.no Particulars Debit($) Credit($)
A Loss on disposal of building 89600
Accumulated depreciation- building(179200*20/40) 89600
Building 179,200
Buildings 354,000
Cash 354,000

C). to record revision in estimated life of the building-if necessary:

No entry required

D). Depreciation for the year 2021(assuming the cost of old roof is removed)

Building =2464000-179200+354000=2,638,800

accumulated depreciation-building (61600*20)-89600=1,142,400
Net=2638800-1142400=1496400
Depreciation for remaining years=1496400/25=59,856

Assume the cost of the new roof is debited to Accumulated Depreciation:
Book value of building prior to replacement=2464000-(61600*20)=1232000
Cost of new roof=336000
Total=1568000
Remaining life=25
Depreciation=1568000/25=62720